THE COST FOR PAINT TO COVER BOTH SIDES OF A LARGE ROAD-SIDE SIGN (FRONT SIDE AND BACK SIDE) WAS $150. A 5-GALLON CAN OF PAINT COSTS $25 AND WILL COVER 250 SQUARE FEET. THE SIGN IS CONSTRUCTED IN THE SHAPE OF A TRAPEZOID. IT HAS ONE BASE MEASURING 35 FT AND ONE BASE MEASURING 40 FT. WHAT IS THE ALTITUDE OF THIS TRAPEZOID?
Answer by checkley71(8403)

ASSUMING IT REQUIRES ALL THE PAINT TO PAINT BOTH SIDES THEN ONE SIDE HAS 750 SQUARE FEET BECAUSE 150/25=6 GALLONS TIMES 250 SQUARE FEET = 1500 SQUARE FEET.
THE AREA OF ONE SIDE IS (35+40)/2 OR 37.5 FEET 37.5H=750 OR H=750/37.5 OR
H-20.